# A Study of Plasticity Loss in On-Policy Deep Reinforcement Learning

### v1

@@ Abstract
Continual learning with deep neural networks presents challenges distinct from both the fixed-dataset and convex continual learning regimes. One such challenge is plasticity loss, wherein a neural network trained in an online fashion displays a degraded ability to fit new tasks. This problem has been extensively studied in both supervised learning and off-policy reinforcement learning (RL), where a number of remedies have been proposed. Still, plasticity loss has received less attention in the on-policy deep RL setting. Here we perform an extensive set of experiments examining plasticity loss and a variety of mitigation methods in on-policy deep RL. We demonstrate that plasticity loss is pervasive under domain shift in this regime, and that a number of methods developed to resolve it in other settings fail, sometimes even resulting in performance that is worse than performing no intervention at all. In contrast, we find that a class of “regenerative” methods are able to consistently mitigate plasticity loss in a variety of contexts, including in gridworld tasks and more challenging environments like Montezuma’s Revenge and ProcGen.
@@ 1 Introduction
In many important machine learning domains, such as continual learning and online reinforcement learning (RL), training data are not wholly available simultaneously. Rather, in these settings, data arrive sequentially over a long period of time. Irrespective of the available quantity of training data, we would typically like to have the highest performing model possible at any given point. Unfortunately, these sequential learning scenarios are known to present optimization issues for neural networks. Specifically, if a model is naively updated to convergence each time new data arrive, and the amount of training data effectively increases, the resulting model generalizes worse than it would under more standard, non-sequential constraints. This phenomenon has been introduced by Ash and Adams, (2020) as the “warm-start problem,” where it was found that randomly initialized models generalize well but are expensive to fit, and warm-started models generalize poorly but converge far more quickly.
An adjacent problem to warm-starting is that of plasticity loss (Lyle et al.,, 2023), which presents itself both in the supervised learning and reinforcement learning settings. Like in the warm-start problem, plasticity loss describes a degradation in a model’s ability to fit new data as training progresses, ultimately harming both the sample efficiency and asymptotic performance. If an agent achieves worse performance when fitting first a source distribution and then a distinct target distribution than it would if instead trained on the target distribution in isolation, the degradation is attributed to plasticity loss.
Whereas the warm-start problem is restricted to test performance, plasticity loss is generally measured as a degradation in performance on training data, often under some form of environmental distribution shift. Despite the typical focus on training data, an ideal solution to plasticity loss would be able to prevent performance degradation both for data seen and unseen when fitting the policy. Importantly, the plasticity loss phenomenon is distinct from overfitting, as resolving plasticity loss should improve both training performance and generalization. In contrast, overfitting describes the opposite, where improving training performance has deleterious effects on generalization.
@@
  • We extend studies of plasticity loss and the warm-start problem to the on-policy regime, finding that they are a persistent issue across a variety of environmental distribution shift conditions.
@@ 2 Preliminaries
An RL algorithm is described as “on policy” if it is trained using data collected from its own policy. While generally more sample inefficient than an off-policy alternative, where the data collection mechanism deviates from the policy that is being trained, on-policy methods have become more commonly used in practical RL problems because of their simplicity and reliability (Andrychowicz et al.,, 2020).
Proximal Policy Optimization (PPO) is a ubiquitous on-policy deep RL algorithm (Schulman et al.,, 2017). As a trust-region method, PPO constrains the extent by which a weight update can modify the current policy, quelling instability issues caused by high-variance policy gradients. Despite this robustness, on-policy algorithms still update on data corresponding only to the agent’s current experience, which may be suboptimal for capturing necessary information around solving the underlying MDP. There is preliminary evidence that plasticity loss can be present in on-policy RL learning problems (Dohare et al., 2023b, ). While prior work has demonstrated this effect in an ad hoc manner, we use multiple environments and distribution change conditions to systematically characterize plasticity loss—and how to resolve it—in the on-policy setting.
@@ 2.1 Simulating environmental distribution shift
Permute. Input pixels are randomly shuffled from round to round. Each of the $k$ environments in a given round are shuffled in the same way, but distinct from the shuffling in any other round. This method is used in Dohare et al., 2023b . We expect that an optimal learning algorithm will be able to achieve equivalent performance on the initial and subsequent rounds.
Window. We supply the agent with $k$ new environments for the same task. Here previously seen environments are no longer accessible, so the agent can only optimize with respect to the current $k$. This approach is considered by Abbas et al., (2023) and Lyle et al., (2023). We expect that an optimal learning algorithm will be able to achieve equivalent or better performance on subsequent rounds as compared to the initial round.
Expand. Like in the window modification, $k$ new environments are supplied to the agent. Here, however, they are appended to the total amount of training data, such that the number of available environments for training at round $n$ is $k\times n$. This is most similar to the warm-start environment studied primarily in Ash and Adams, (2020) and revisited in Igl et al., (2020). In this setting we expect that even for an optimal learning algorithm training performance to decrease each round, as the dataset becomes more difficult to fit. Correspondingly, we expect the generalization performance to increase as the training distribution expands to better capture the true distribution.
In the case of the latter two modifications, data at each new round are drawn from the exact same distribution as environments from previous rounds. As we demonstrate in Section 4.1, these still induce significant performance degradation in warm-started models. This observation is surprising—we often think of a model’s initialization as a sort of “prior” over learned functions, and here we demonstrate that even an initialization obtained by training on data distributed identically to the current round still hinders performance in on-policy learning.
@@ 3 Existing Approaches
A number of methods have been proposed to address plasticity loss in deep neural networks, though none were designed with on-policy reinforcement learning explicitly in mind. Instead, these methods have been primarily validated in the context of either supervised learning or off-policy RL. Each of these techniques can be divided into one of two groups: Interventions which are performed intermittently during training, and those which are applied continuously, either as part of the model architecture or as part of the model update process. For additional implementation details see Appendix Section A.
@@ 3.1 Intermittent Interventions
We consider an intervention to be intermittent if it is applied only at specific points during training. With one exception, each time the training data change, the interventions in this category are applied and training proceeds. Note that this implies that there must be awareness of when the training distribution shifts, which in practice may be difficult or infeasible.
Resetting final layer. Proposed by Nikishin et al.,, this method involves periodically replacing the weights of the final layer in the network with newly initialized values. It was demonstrated that this alleviates plasticity loss in certain off-policy RL.
Shrink+Perturb. This technique periodically reduces the magnitude of all weights in the network by a factor and then adding noise sampled from a new weight initialization which is scaled down by a separate factor (Ash and Adams,, 2020). It was demonstrated that this improves performance in batched continual learning settings. In our implementation the weight of these two factors are entangled such that they sum to one.
Plasticity Injection. Plasticity Injection replaces the final layer of a network with a new function that is a sum of the final layer’s output and the output of a newly initialized layer subtracted by itself (Nikishin et al.,, 2023). The gradient is then blocked in both the original layer and the subtracted new layer. It was shown that this method increases performance of off-policy RL agents in the Atari Learning Environment (ALE).
ReDo. This technique resets individual neurons withing the network based on a dormancy criteria at fixed intervals (Sokar et al.,, 2023), leading to improvement in performance in the context of off-policy RL agents trained on the ALE. Following Sokar et al.,, we reset the model parameters more frequently than we apply environmental distribution shifts.
@@ 3.2 Continuous Interventions
LayerNorm. LayerNorm (Ba et al.,, 2016) is a ubiquitous deep learning regularization technique, and was shown useful for plasticity loss by Lyle et al.,. The authors demonstrated a performance improvement in the ALE when using off-policy RL.
CReLU activations. Proposed by Abbas et al., and studied further by Lee et al.,, this method involves replacing ReLU activations with the CReLU activation function (a concatenation of $\mathrm{ReLU}(x)$ and $\mathrm{ReLU}(-x)$ for an input scalar $x$). This ensures that the gradient is non-zero for all units in a given layer, which appears to mitigate plasticity loss when periodically switching between tasks in the ALE with off-policy methods.
Regenerative Regularization. Regenerative regularization uses an L2 penalty in the optimized loss to encourage model weights to be near their initial values (Kumar et al.,, 2023). It has been shown to be effective in a number of continual learning settings.
@@ 4 Experiments
Figure 2: Performance in the gridworld environment under each of the three distribution shift conditions. The degradation between rounds is evidence of plasticity loss. (A): Epoch-level training performance for permute modification. Dotted vertical lines indicate the end of each round, before a new environmental distribution shift is applied. (B): Round-level training performance for the Permute modification. Data points correspond to normalized mean reward in final 50 episodes of round. (C, D): Round-level training performance for the Window and Expand conditions. Top row: Training performance. Bottom row: Test performance.
Below we present results of a suite of experiments conducted using a 2D gridworld environment (Juliani et al.,, 2022), the procedurally generated CoinRun environment (Cobbe et al.,, 2019), and the Atari game Montezuma’s Revenge (Bellemare et al.,, 2013). All models are trained using PPO (Schulman et al.,, 2017). For the gridworld experiments we use an MLP encoder, and for the CoinRun and Montezuma’s Revenge experiments we instead use a convolutional encoder. We set the number of environment instances ($k$) to 100 for all experiments conducted with the gridworld and CoinRun environments. Like in previous sections, many of these results show reward trends as a function of the number of rounds, with each round denoting a new shift in the state distribution. These curves are shown in terms of normalized reward (i.e. we subtract the reward obtained after the first round from the reward obtained after each subsequent round) to compensate for the fact that several algorithms have a regularization effect that improves performance on average but does not necessarily improve the performance trend—it is possible to improve generalization, for example, without remedying the degradation in train or test performance induced by plasticity loss. Training details can be found in Appendix Section B.
@@ 4.1 Plasticity loss is apparent in on-policy RL
We find evidence of plasticity loss in the on-policy reinforcement learning setting under multiple kinds of distribution shift. Figure 2 presents these results. As mentioned earlier, for round-level plots like these, reward values are normalized such that the mean reward at the end of the first round is set to zero. As such, positive and negative values correspond respectively to increases and decreases in performance relative to the first round. Decreasing normalized reward indicates a decay in relative performance and is indicative of plasticity loss.
For both the permute and window shift conditions, Figure 2 demonstrates clear degradation in mean episodic reward for the warm-started model as compared to a model which has all of its weights reset between each round. While there is some positive transfer for the warm-start method in the expand shift condition, we find generalization performance suffers here in comparison to reset-all. We provide statistical tests comparing these methods to warm-start and random initialization baselines in Appendix Section D.
@@ 4.2 Predictors of plasticity loss and generalization
Figure 3: Top: Correlation plots of normalized mean reward in the gridworld environment compared against various model attributes. Each point is averaged over five replicates, and shows the final values produced after a ten-round experiment. Values for each measurement are normalized by its baseline level at initialization. Measurements that significantly correlate ($p<0.05$) with normalized reward are bolded. First row: Training performance. Second row: Test performance. Bottom: Values of the three predictive attributes during the course of training for each intervention and window change condition as compared to training performance. Final values of plots like these correspond to a single point in the correlation plots in the first two rows.
@@
Identifying the exact statistical underpinnings of plasticity loss is an active area of research in continual and reinforcement learning. Here we identify several factors which are significantly correlated with plasticity loss as well as a degradation in generalization performance. Figure 3 presents correlation plots of normalized reward with various attributes of interest.
@@
Overall, we find that weight magnitude and number of dead units are both significantly correlated with the normalized reward [$p<0.05$]. Here, “dead units” refers to ReLU activations that never produce non-negative outputs. Note that the LayerNorm and CReLU methods are not included in this analysis, due to their immediate impact on the weight magnitude and dead unit counts, respectively. For plots of these measurements over time during training, see Supplemental Figure 9.
@@
We find that neither the gradient magnitude nor the magnitude of weight difference between updates has a significant relationship with either plasticity or generalization. This is in contrast to the findings of Abbas et al., and Nikishin et al.,, each of which examined plasticity loss in the off-policy rather than on-policy setting. The former found that weight difference between SGD updates is correlated with loss of plasticity while the latter found that gradient norms were predictive.
@@
We also fit a generalized linear model (GLM) using all five criteria as independent variables and normalized reward as the dependent variable. Here we find that both gradient and weight magnitudes are significantly predictive of train performance [$p<0.01$], while dead unit count no longer is [$p>0.05$]. This suggests that the predictive power of the dead unit count can be accounted for by the weight magnitude. An additional GLM was fit to the test data, and we find that only the dead unit count remains predictive [$p<0.05$].
@@
Although these analyses do not establish causality, the robust significance of weight magnitude in predicting plasticity loss suggests that plasticity loss may be a function of how much trainable parameters deviate from their initialization distribution. As such, successful mitigation strategies might curb unbounded increases in the weight norm of network parameters over the course of learning. This has the downstream effect of minimizing the number of dead units in the network. Below we compare mitigation strategies, finding that those that keep weight norm from excessively growing seem to perform best.
4.3 Architectural interventions do not fully address plasticity loss

### v2

@@ Abstract
Continual learning with deep neural networks presents challenges distinct from both the fixed-dataset and convex continual learning regimes. One such challenge is plasticity loss, wherein a neural network trained in an online fashion displays a degraded ability to fit new tasks. This problem has been extensively studied in both supervised learning and off-policy reinforcement learning (RL), where a number of remedies have been proposed. Still, plasticity loss has received less attention in the on-policy deep RL setting. Here we perform an extensive set of experiments examining plasticity loss and a variety of mitigation methods in on-policy deep RL. We demonstrate that plasticity loss is pervasive under domain shift in this regime, and that a number of methods developed to resolve it in other settings fail, sometimes even performing worse than applying no intervention at all. In contrast, we find that a class of “regenerative” methods are able to consistently mitigate plasticity loss in a variety of contexts, including in gridworld tasks and more challenging environments like Montezuma’s Revenge and ProcGen.
@@ 1 Introduction
In many important machine learning domains, such as continual learning and online reinforcement learning (RL), training data are not wholly available simultaneously. Rather, in these settings, data arrives sequentially over a long period of time. Irrespective of the available quantity of training data, we would typically like to have the highest performing model possible at any given point. Unfortunately, these sequential learning scenarios are known to present optimization issues for neural networks. Specifically, if a model is naively updated to convergence each time new data arrives, and the amount of training data effectively increases, the resulting model generalizes worse than it would under more standard, non-sequential constraints. This phenomenon has been introduced by Ash and Adams, (2020) as the “warm-start problem,” where it was found that randomly initialized models generalize well but are expensive to fit, and warm-started models generalize poorly but converge far more quickly.
An adjacent problem to warm-starting is that of plasticity loss (Lyle et al.,, 2023), which presents itself both in the supervised learning and reinforcement learning settings. Like in the warm-start problem, plasticity loss describes a degradation in a model’s ability to fit new data as training progresses, ultimately harming both sample efficiency and asymptotic performance. If an agent achieves worse performance when fitting first a source distribution and then a distinct target distribution than it would if instead trained on the target distribution in isolation, the degradation is attributed to plasticity loss.
Whereas the warm-start problem is restricted to test performance, plasticity loss is generally measured as a degradation in performance on training data, often under some form of distribution shift. Despite the typical focus on training data, an ideal solution to plasticity loss would be able to prevent performance degradation both for data seen and unseen when fitting the policy. Importantly, the plasticity loss phenomenon is distinct from overfitting, as resolving plasticity loss should improve both training performance and generalization. In contrast, overfitting describes the opposite, where improving training performance has deleterious effects on generalization.
@@
  • We extend studies of plasticity loss and the warm-start problem to the on-policy regime, finding that they present a persistent issue across a variety of model architectures and environmental distribution shift conditions.
@@ 2 Preliminaries
An RL algorithm is described as “on policy” if it is trained using data collected from its own policy. While generally more sample inefficient than off-policy alternatives, where the data collection mechanism deviates from the policy being fit, on-policy methods have become more commonly used in practical RL problems because of their simplicity and stability (Andrychowicz et al.,, 2020).
Proximal Policy Optimization (PPO) is a ubiquitous on-policy deep RL algorithm (Schulman et al.,, 2017). As a trust-region method, PPO constrains the extent by which a weight update can modify the current policy, quelling instability issues caused by high-variance policy gradients. Despite this robustness, on-policy algorithms still update on data corresponding only to the agent’s current experience, which may be sub-optimal for capturing necessary information around solving the underlying MDP. There is some evidence that plasticity loss can be present in on-policy RL learning problems (Dohare et al., 2023b, ), demonstrating the effect in a somewhat ad hoc manner, we systematically use multiple environments and distribution shift conditions to more thoroughly characterize plasticity loss—and how to resolve it—in the on-policy setting.
@@ 2.1 Simulating environmental distribution shift
Permute. Input pixels are randomly shuffled from round to round. Each of the $k$ environments in a given round are shuffled in the same way, but distinct from the shuffling in any other round. This method was utilized in Dohare et al., 2023b . We expect that an optimal learning algorithm will be able to achieve equivalent performance on the initial and subsequent rounds.
Window. We supply the agent with $k$ new environments for the same task. Here previously seen environments are no longer accessible, so the agent can only optimize with respect to the current $k$. This approach was considered by
Abbas et al., (2023) and Lyle et al., (2023). We expect that an optimal learning algorithm will be able to achieve equivalent or better performance on subsequent rounds as compared to the initial round.
Expand. Like in the window modification, $k$ new environments are supplied to the agent. Here, however, they are appended to the total amount of training data, such that the number of available environments in the final round of training is $k\times n$. This is most similar to the warm-start environment studied primarily in Ash and Adams, (2020), and revisited in Igl et al., (2020). In this setting we expect training performance to decrease each round, even for an optimal learning algorithm, as available training data grows and becomes more difficult to fit. Correspondingly, we expect the generalization performance to increase as the training distribution expands to better capture the true distribution.
In the case of the latter two modifications, data at each new round are drawn from the exact same distribution as environments from previous rounds. As we demonstrate in Section 4.1, these still induce significant performance depredation in warm-started models. This observation is surprising—we often think of a model’s initialization as a sort of “prior” over learned functions, and here we demonstrate that even an initialization obtained by training on data distributed identically to the current round still hinders performance in on-policy learning.
@@ 3 Existing Approaches
A number of methods have been proposed to address plasticity loss in deep neural networks, though none were designed with on-policy RL explicitly in mind. Instead, these methods have been primarily validated in the context of either supervised learning or off-policy RL. Each of these techniques can be divided into one of two groups: Interventions which are performed intermittently during training, and those which are applied continuously, either as part of the model architecture or as part of the model update process. For additional implementation details see Appendix Section A.
@@ 3.1 Intermittent Interventions
We consider an intervention to be intermittent if it is applied only at specific points during training. Periodically, the interventions in this category are applied, and training proceeds normally otherwise. Most of these are applied each time there is a training distribution change, implying that there must be awareness of when this occurs; in practice this information may be unavailable and difficult to detect.
Resetting final layer. Proposed by Nikishin et al.,, this method involves periodically replacing the weights of the final layer in the network with newly initialized values. It was demonstrated that this alleviates plasticity loss in certain off-policy RL settings.
Shrink+Perturb. This technique periodically scales the magnitude of all weights in the network by a factor and then adds a small amount of noise (Ash and Adams,, 2020). It was demonstrated that this improves performance in batched continual learning settings. In our implementation the weight of these two factors are entangled such that they sum to one.
Plasticity Injection. Plasticity Injection replaces the final layer of a network with a new function that is a sum of the final layer’s output and the output of a newly initialized layer subtracted by itself (Nikishin et al.,, 2023). The gradient is then blocked in both the original layer and the subtracted new layer. It was shown that this method increases performance of off-policy RL agents in the ALE.
ReDo. This technique resets individual neurons within the network based on a dormancy criteria at fixed intervals (Sokar et al.,, 2023), leading to improvement in performance in the context of off-policy RL agents trained on the ALE. Following Sokar et al.,, we reset the model parameters more frequently than we apply environmental distribution shifts.
@@ 3.2 Continuous Interventions
LayerNorm. LayerNorm (Ba et al.,, 2016) is a ubiquitous deep learning regularization technique, and was shown useful for mitigating plasticity loss by Lyle et al.,. The authors demonstrated a performance improvement in the ALE when using off-policy RL.
CReLU activations. Proposed by Abbas et al., and studied further by Lee et al.,, this method involves replacing ReLU activations with the CReLU activation function. This ensures that the gradient is non-zero for all units in a given layer, which appears to mitigate plasticity loss when periodically switching between tasks in the ALE with off-policy methods.
Regenerative Regularization. Regenerative regularization uses an L2 penalty to encourage model weights to be near their initial values (Kumar et al.,, 2023). It has been shown to be effective in a number of continual learning settings.
@@ 4 Experiments
Figure 2: Performance in the gridworld environment under each of the three distribution shift conditions. The degradation between rounds is evidence of plasticity loss. (A): Epoch-level training performance for permute modification. Dotted vertical lines indicate the end of each round, before a new environmental distribution shift is applied. (B): Round-level training performance for the Permute modification. Data points correspond to normalized mean reward in final 50 episodes of round. Shaded regions correspond to standard error. (C, D): Round-level training performance for the Window and Expand conditions. Top row: Training performance. Bottom row: Test performance.
Below we present results of a suite of experiments conducted using a 2D gridworld environment (Juliani et al.,, 2022), the procedurally generated CoinRun environment (Cobbe et al.,, 2019), and the Atari game Montezuma’s Revenge (Bellemare et al.,, 2013). All models are trained using PPO (Schulman et al.,, 2017). For the gridworld experiments we use an MLP encoder, and for the CoinRun and Montezuma’s Revenge experiments we instead use a convolutional encoder. We set the number of environment instances ($k$) to 100 for all experiments conducted with the gridworld and CoinRun environments. Additional training details can be found in Appendix Section B.
@@ 4.1 Plasticity loss is apparent in on-policy RL
We find evidence of plasticity loss in the on-policy RL setting in the presence of multiple kinds of distribution shift, as presented in Figure 2. For round-level plots like these, reward values are normalized such that the mean reward at the end of the first round is set to zero. As such, positive and negative values correspond respectively to increases and decreases in performance relative to the first round. Decreasing normalized reward indicates a decay in relative performance, implying plasticity loss.
For both the permute and window shift conditions, Figure 2 demonstrates clear degradation in mean episodic reward for the warm-started model as compared to a model which has all of its weights reset between each round. While there is some positive transfer for the warm-start method in the expand shift condition, we find generalization performance suffers in this condition as compared to reset-all. We provide statistical tests comparing the performance of these methods to the warm-start and random initialization baselines in Section D of the appendix.
@@ 4.2 Predictors of plasticity loss and generalization
Figure 3: Top: Correlation plots of normalized mean reward in the gridworld environment compared against identified metrics. Each point is averaged over five replicates, and shows the final values produced after a ten-round experiment. Values for each measurement are normalized by its baseline level at initialization. Measurements that significantly correlate ($p<0.05$) with normalized reward are bolded. First row: Training distribution performance. Second row: Test distribution performance. Bottom: Values of the three predictive metrics during the course of training for each intervention and window change condition as compared to training performance. Shaded regions correspond to standard error. Final values of plots like these correspond to a single point in the correlation plots above.
@@
Identifying the exact statistical underpinnings of plasticity loss is an active area of research in continual and reinforcement learning. Here we identify several factors which are significantly correlated with plasticity loss as well as a degradation in generalization performance. Figure 3 presents correlation plots of normalized reward with various metrics of interest. Here we normalize rewards (i.e. subtract the initial reward value from the final reward value) to compensate for the fact that many of these algorithms have a regularization effect that improves performance on average but does not necessarily improve performance trends—it is possible to improve generalization, for example, without remedying plasticity loss.
@@
Note that the LayerNorm and CReLU methods are not included in this analysis, due to their direct effects on the weight magnitude and dead unit counts, respectively. For graphs of the underlying metrics over time during training, see Supplemental Figure 11. We find that weight magnitude and number of dead units are both significantly correlated with the normalized reward [$p<0.05$]. “Dead units” refer to ReLU activations that never produce non-negative outputs.
@@
We find that neither the gradient magnitude nor the magnitude of weight difference between updates has a significant, linear relationship with either plasticity or generalization. This is in contrast to the findings of Abbas et al., and Nikishin et al.,, each of which examined plasticity loss in off-policy rather on-policy settings. The former found that weight difference was correlated with loss of plasticity while the latter found that gradient norms were predictive.
@@
We also fit a generalized linear model (GLM) using all five metrics as independent variables and normalized reward as the dependent variable. Here we find that both gradient and weight magnitudes are significantly predictive of train performance [$p<0.01$], while dead unit count no longer is [$p>0.05$]. This suggests that the predictive power of the dead unit count can be accounted for by the weight magnitude. An additional GLM was fit to the test data, and we find that only the dead unit count remains predictive [$p<0.05$].
@@
Although these analyses do not establish causality, the robust significance of weight magnitude in predicting plasticity loss suggests that plasticity loss may be a function of how much trainable parameters deviate from their initialization distribution. As such, successful mitigation strategies would induce a sub-linear increase in the weight norm of the network parameters over the course of learning. This has the downstream effect of minimizing the number of dead units in the network. Below we compare the mitigation strategies, finding that those which address this underlying cause directly indeed perform best.
4.3 Novel architectural methods do not fully address plasticity loss